OpenStudy (anonymous):

I need help with Volume and Radius! Betelgeuse is one of the stars found in the constellation orion. its radius is about 1500 times the radius of the sun. How many times as great as the sun's volume is Betelgeuse's volume?

OpenStudy (anonymous):

Betelgeus r = 1500(s) where s = the suns radius assuming the sun is a sphere (I don't think it is but I could be wrong) we acquire the volume with the formula V = 4πr^(3)/3 so V = 4π(1500(s))^(3)/3 To solve this we can just assign an arbitrary number to s because the radius of Betelgeus is consistently 1500 greater and thus the volume will be consistently larger by a specific amount. So I'm just going to say the suns radius is 1 this s = 1 V = 4π(1500(1))^(3)/3 V = 4 500 000 000 Now lets check the suns Area based on it being 1 V = 4π((1))^(3)/3 Thus its Volume is equal to 4π((1))^(3)/3 Now just divide the numbers to see how many times the volume of the sun can go into the volume of betelgeus 4 500 000 000/4π((1))^(3)/3 = 3,375,000,000/pi x greater I could be wrong though

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

am I wrong or do you know lol

OpenStudy (anonymous):

i dont know... i dont have the answer

OpenStudy (anonymous):

fair enough I think my logic is pretty sound in this one though but meh

OpenStudy (anonymous):

ok